Circles has radius 5 and tangent to the line 3x+4y=24 at the point (2, 412⁄). Illustrate thenfindthe equations of thecircles


1
Expert's answer
2022-02-03T08:53:14-0500

Solution:

"r=5"

Circle is tangent to the line 3x+4y=24 at (2, 41/2)

Or tangent is 3x+4y=24.

Slope of tangent "=-\dfrac 34"

Let the equation of the circle be "(x-h)^2+(y-k)^2=r^2"

"\Rightarrow (x-h)^2+(y-k)^2=5^2
\\ \Rightarrow (x-h)^2+(y-k)^2=25\ ...(i)"

On differentiating w.r.t "x",

"2(x-h)+2(y-k)y'=0
\\\Rightarrow y'=\dfrac{x-h}{y-k}\ \ \ \ \ [Slope\ of\ tangent]"

For (2, 41/2), put "x=2, y=41\/2, y'=-3\/4" in above equation.

"-\dfrac 34=\dfrac{2-h}{\frac {41}2-k}
\\ \Rightarrow h=-\dfrac{4k-88}{3}"

There is no other info in the problem, as the question is incomplete.

Assume h=0, then k=22

Now put it in (i), we will get the equation of circle.

"x^2+(y-22)^2=25"